# Customizable wave tailoring materials enabled by coupling nonlinear inverse design at two scales

### v1

@@ Abstract
Passive transformation of waves via nonlinear systems is ubiquitous in settings ranging from acoustics to optics and electromagnetics. Passivity is of particular importance for responding rapidly to stimuli, and nonlinearity enormously expands signal transformability compared to linear systems due to the breaking of superposition. It is well known that different types of nonlinearity yield vastly and qualitatively different effects on propagating signals, which raises the question of “what precise nonlinearity is the best for a given wave tailoring application?” This question has largely remained in the regime of simulation and theory, as, until recently it has not been possible to freely choose any optimal nonlinearity (instead relying on limited tunability around known nonlinear mechanisms). Herein, we leverage recent advances in mechanics, wherein desired nonlinear constitutive responses can be achieved through shape optimization, and couple this to a larger-scale, reduced-order nonlinear dynamical inverse design step. Using minimization of kinetic energy transmission from impact as a case study, we identify ideal nonlinear constitutive responses and the geometries needed to achieve them within a single design process. As part of this, we show the large sensitivity of this metric to small changes in nonlinearity, and thus the need for high precision, free-form nonlinearity tailoring. We validate our predictions using impact experiments in a chain of nonlinear springs and masses. This work sets the foundation for broader passive nonlinear mechanical wave tailoring material design, including future incorporation of higher-dimensionality, irreversibility, material heterogenity, and coupled physics. Potential applications include mechanical computing, acoustic signal and image processing, impact and vibration mitigation, and materials that autonomously respond and conduct directed work via rapid shape-change.
keywords:
inverse design, nonlinear dynamics, nonlinear mechanics, impact mitigation, nonlinear waves
1 Introduction
@@
In this work, we introduce a method to create customizable wave tailoring materials by coupling nonlinear response inverse design at two scales. Namely, we optimize for the emergent dynamic response of the bulk system using a reduced order, discrete element model (DEM) simulation to identify an optimal nonlinear constitutive law for the given performance metric, and couple this to a unit-cell-scale, geometrically-nonlinear, free-form, shape optimization algorithm which designs a physical system that achieves the nonlinear constitutive property identified by the DEM (outlined in Fig. 1). Further, we demonstrate the clear need for high precision nonlinear response design, wherein the system dynamic performance is shown to be highly sensitive to the nonlinear coefficients of the constitutive law. Herein, we choose the maximum kinetic energy transmitted to the boundary of the system in response to an impact event, normalized by that of an otherwise-identical linear system, as our performance metric. The comparison between the linear and nonlinear response is particularly important, as it isolates the role of nonlinearity from other wave manipulating effects such as dispersion and dissipation. We demonstrate the full inverse design of this superior impact mitigation system, from identification of an ideal nonlinearity, to the design of the unit cell geometry, and experimental validation of the performance. An important feature to note, unlike some prior computational quasi-static nonlinear mechanical design strategies [22, 37, 24, 25], we do not use simplified or reduced order models for our underlying mesostructure design, which enables a broader design space and access to highly precise tailoring of nonlinear responses [23]. While we demonstrate our two-scale inverse design for the case of impact mitigation, we expect such an approach for the previously specified nonlinear wave transformation applications [1, 7, 8, 9, 10, 11, 12, 13] within the acoustic, phononic, and mechanical wave settings.
@@
Figure 1: Overview of the multilevel design flow. (A) Discrete element model (DEM) simulation of the entire system dynamics. (B) Identification of the optimal nonlinear constitutive law. (C) Shape optimization of a mesostructure to match the identified nonlinear constitutive law. (D) Mechanical and dynamic experimental system characterization.
2 Results
2.1 Reduced order discrete element model
@@
The DEM (Fig. 1A) simulates the impact of a rigid “impactor” block onto a lattice (chain) with $N$ unit cells and $a$ unit cell length. We represent the chain as lumped masses of mass $m$, interconnected by massless nonlinear springs and inter-site linear dampers, emulating the behavior of a viscoelastic material. The normalized force-displacement relationship for a unit cell of the material is modeled using an up-to third-order polynomial (denoted in Fig. 1B), expressed as $f(x)=x+c_{2}x^{2}+c_{3}x^{3}$ (where the force is normalized the linear stiffness $c_{1}^{\ast}$ and $x<0$ denotes compression, with $x$ the extension of the spring normalized by $a$). In tension, the response is set to that of a linear spring, such that $f(x)=x$. Incorporating the nonlinear spring into dimensionless equations of motion of the chain we obtain
@@
where $\zeta$ is the inter-site damping ratio and $x_{i}$ is the displacement of the $i$th particle. A variable mass and velocity rigid impactor is used on the left, and a fixed boundary is used on the right. The model also incorporates a contact spring designed to facilitate the smooth contact and controlled release of the impactor during initial impact and rebound, respectively. The simulated dynamical response is acquired through the numerical integration of Eq. (1) via a Runge-Kutta algorithm. The non-dimensionalization of all variables and full details concerning the equations of motion are described in SI Note 1.
2.2 Optimization of nonlinear constitutive law based on dynamical response
@@
Before searching for optimal nonlinear constitutive responses with our DEM, we set several bounds. First, we confine the unit cell strain to $1$ in compression, and set $c_{3}>0$ for simplicity. Second, we restrict our search range for nonlinear coefficients $c_{2}$ and $c_{3}$ to ensure positive strain energy throughout the entire compression range. By examining the polynomial’s properties within this range, we classify the quasi-static response into three distinct zones, “bistability”, “monotonic increase”, and “local maximum”, as shown in Fig. 2A. Bistability (magenta area) denotes the existence of both a local maximum and minimum other than the boundaries (the local minimum does not need to fall below zero). Monotonic increase (blue area) denotes the absence of extrema. Local maximum (green area) signifies the presence of a local maximum (no local minimum existed) within the range of the length of one unit cell (more details in SI Note 2).
@@
Figure 2: Identification of optimal nonlinear constitutive response via DEM simulation for a single impact condition. (A) Feasible solutions of nonlinear spring coefficients $c_{2}$ and $c_{3}$. The black line represents $c_{2}=-\sqrt{3c_{3}}$, the red line indicates $c_{2}=-(1+3c_{3})/2$, the blue line is $c_{2}=-3c_{3}$, and the green line is the zero strain energy throughout the whole range, $c_{2}=-3/2-3c_{3}/4$. (B) Non-dimensional force displacement relationship of the best performance of the nonlinear spring. (C) Ratio of maximum kinetic energy at the last particle in the nonlinear chain to the initial impact energy as a function of nonlinear spring coefficients for the impact condition of $M/M_{0}=0.05$ and $V/V_{0}=1$, as well as $\zeta=0.01$. The lines from (A) are overlaid, and the star marker denotes the point of best performance. Normalized kinetic energy of the (D) best performing nonlinear and (E) linear material.
@@
Within these bounds, we first optimize for a singular impact condition ($M/M_{0}=0.05$ and $V/V_{0}=1$), where $M_{0}$ is half the mass of the chain and $V_{0}$ is the linear sound speed. The material is composed of $20$ particles and $\zeta=0.01$. We vary the nonlinear coefficients of the springs $c_{2}$ and $c_{3}$, aiming to minimize the maximum kinetic energy experienced at the end of the material ($KE_{non}$) normalized by that of a linear system ($KE_{lin}$) which has all of the same properties except $c_{2}=c_{3}=0$. Figure 2C quantifies the ratio of maximum kinetic energy at the end of the material to the initial impact energy ($TE$), as a function of the coefficients of the nonlinear spring. Around the region of best performance in Fig. 2C ($f(x)=x+5.81x^{2}+7.08x^{3}$, plotted in Fig. 2B, and denoted by the star marker in the inset of Fig. 2C) significant performance sensitivity to the constituent nonlinear parameters can be seen (see the zoomed-in view in Fig. 2C). To describe this sensitivity more quantitatively, in SI Note 3 we calculate the gradient of the inset in Fig. 2C, which shows $\partial(KE_{non}/TE)/\partial c_{2}$ and $\partial(KE_{non}/TE)/\partial c_{2}$ approaching ±$\pm 2$. Spatiotemporal responses of kinetic energy of the optimal nonlinear and linear chain are shown in Fig.  2(D) and (E), respectively. While both materials exhibit an initially sharp pulse, energy from the optimal nonlinear material is trapped around particle 15, preventing further transmission.
@@
In the subsequent analysis, we conduct an optimization wherein we look for optimal nonlinear coefficients for varied impactor mass and velocities. We use the same ratio of the maximum $KE$ of the last particle of the nonlinear material normalized by that of the linear material as a performance metric (denoted henceforth as “$KE$ ratio”), wherein a lower ratio indicates superior performance of the nonlinear material. The optimal (minimum) $KE$ ratios with respect to $M/M_{0}$ and $V/V_{0}$ and corresponding nonlinear spring parameters $c_{2}$ and $c_{3}$ can be seen in Fig. 3. In contrast to Fig. 2, we use a lower damping ($\zeta=0.005$), chosen to emulate that of the polycarbonate springs used in our experimental realization. Additional simulation results of $KE$ ratios for increased damping and greater discreteness (more unit cells) are available in SI Note 4 and indicate the potential for $KE\text{ratio}<10^{-2}$ in the latter case. The damping value used in Fig. 3 was chosen to approximate damping found by measuring the resonance of a single connector and nonlinear spring unit (see SI Note 5). Simulation cases of the nonlinear material where self-contact occurs are discarded from consideration of the optimal performance.
@@
Figure 3: Identification of optimal nonlinear coefficients via DEM simulation for varied impactor conditions. (A) Optimal kinetic energy ratio ($KE$ ratio) as a function of impact conditions and corresponding nonlinear spring coefficients (B) $c_{2}$ and (C) $c_{3}$ for a material of $N=20$ and $\zeta=0.005$. The color bar in (A) is saturated at a $KE$ ratio of unity. The dashed black line denotes the onset of self contact ($x<-1$) in the linear system. The dot in (A) denotes the high performing case used for mesostructure design.
@@
There is a clear boundary where the nonlinear chain does not outperform the linear, which is correlated with the occurrence of self-contact within the linear chain (dashed black line in Fig. 3). At impactor velocities and masses below this threshold, the nonlinear materials exhibit significantly enhanced mitigation effectiveness. As can be seen in Fig. 3A, at a relatively low to moderate level of impactor mass ($\sim 0.01$ to $0.2M_{0}$), optimal nonlinear materials exhibit a mitigation capability ($KE$ ratio) that is over ten times better than the linear material. As the impactor mass and velocity increase and cross over the dashed black line (orange to red area near the dashed line in Fig. 3A), the priority shifts to preventing contact between unit cells, leading to a comparatively impaired energy-absorbing performance. In future studies, such self contact could be explored as a form of nonlinearity and design feature, instead of a constraint.
@@
A specific optimal solution (pair of coefficient ratios) was chosen from a set of multiple solutions showing high performance $KE$ ratio. We further downselected, looking for a solution also with relatively low strain, making it more amenable to experimental implementation. The chosen solution is a nondimensional nonlinear mechanical response of the form $f(x)=x+94.4x^{2}+2000x^{3}$ in compression, with an impact condition of $M/M_{0}=0.0148$ and $V/V_{0}=0.11$ (marked in Fig. 3A). This nonlinear constitutive law is used as the target for the mesostructure design, with the goal of finding a geometry that yields that nonlinear mechanical response.
2.3 Shape optimization for desired nonlinear constitutive law
@@
In order to find a spring geometry that gives the desired nonlinear response, a two-dimensional shape optimization approach is taken, using a level-set optimization method [23]. A third order polynomial is fit to the calculated force-displacement behavior of the structure, and the objective of the optimization problem is taken as the ratio of nonlinear to linear terms. The exact form of the objective function is
@@
where $c_{i}$ and $c_{j}$ represent the current polynomial coefficients, and $T_{i}$ and $T_{j}$ represent the target coefficients. By taking the ratio of the polynomial terms, the nonlinearity of the structure is decoupled from the linear stiffness, allowing the optimizer more design freedom, leading to more robust convergence. To further aid in the navigation of the design space, we select an initial condition for our optimization process which displays qualitatively similar behavior to that desired (see SI note 5).
@@
When realizing our nonlinear chain we employ comparatively rigid frames around the designed spring to mimic fixed boundary conditions (as can be seen in Fig. 1D) and rigid connectors between the springs to allow relative movement of the particles. With the addition of these components we have a unit cell length, $a$, and a nonlinear spring design domain length that encompasses only a portion of this larger unit cell. We call this portion of $a$ the spring length, and denote it by $a_{s}$ (and, hereafter, the subscript $s$ is used to refer to parameters defined on the scale of the spring). Figure 4E highlights the difference between these two length scales. Because the DEM-identified polynomial constitutive law is expressed as a function of strain, and the strain experienced by the spring across $a_{s}$ is different than that experienced by the entire unit cell across $a$ for a fixed applied displacement, the targeted polynomial is therefore scaled accordingly. We take the ratio of nonlinear terms as $R=(c_{i}\epsilon^{i})/(c_{1}\epsilon)$ and $R_{s}=(c_{s,i}\epsilon_{s}^{i})/(c_{s,1}\epsilon_{s})$ on the scales of $a$ and $a_{s}$, respectively, with $i$ representing the order (or power) of the term and $\epsilon$ representing the maximum strain experienced on the corresponding length scale. In order for equivalent degrees of nonlinearity to be displayed at different scales, we take $R=R_{s}$ at the maximum strain considered in each scale, and solve for updated $c_{i}$ or $c_{s,i}$ terms.
@@
The optimization target (recall, on the unit cell scale, identified above as $f(x)=x+94.4x^{2}+2000x^{3}$) can therefore be expressed on the spring length scale as $f=x_{s}-30.2x_{s}^{2}+204.8x_{s}^{3}$, by setting $R=R_{s}$ and solving for $c_{s,i}$. The final optimized structure, shown in Fig. 4A-B, achieved an experimental force-displacement law on the spring scale of $f(x)=-0.0003+x_{s}-28.28x_{s}^{2}+193.13x_{s}^{3}$, resulting in a percent difference between targeted and obtained polynomial ratios of $5.70\%$ for the third order ratio and $6.34\%$ for the second order ratio.
@@
Figure 4: Shape optimization to identify geometry giving DEM identified desired constitutive law, and experimental realization of the spring and chain. A-B) Results of the shape optimization, with the target and final force-displacement curves shown in (A), and the optimized design shown in (B). C) The fabricated polycarbonate unit cell, consisting of four optimized springs (as seen in (B) and a rigid frame to impose boundary conditions. D) The quasi-static test of the unit cell shown in (C), compared against the behavior of a single spring with perfectly imposed boundary conditions as simulated via Comsol FEM simulation. E) The full chain of 20 unit cells, hung from a frame. The chain is clamped to the left of the leftmost unit cell, imposing a zero displacement boundary condition. The impact occurs at the right end of the chain. The unit cell and spring length scales are labeled.
@@
As the multilevel optimization methodology is a loop in which inputs from the DEM and shape optimization feed back into one another, the design process can be followed in either direction, which can be exploited for exploration of the space. For example, while one may begin with a desired impact condition which is then fed into the DEM simulation followed by the mesostructure optimization, one might just as easily begin with some nonlinearity and feed this into the DEM simulation to explore what impact conditions it might perform well (or poorly) for. Similarly, one might find that the physical design domain requires constraints, or that optimized structures exhibit structural damping which diverges significantly from predictions, and these updated considerations can be flowed back up to the DEM simulation to search for the updated optimum conditions.

### v2

# Customizable wave tailoring materials enabled by nonlinear bilevel inverse design

@@ Abstract
Passive transformation of waves via nonlinear systems is ubiquitous in settings ranging from acoustics to optics and electromagnetics. Passivity is of particular importance for responding rapidly to stimuli and nonlinearity enormously expands signal transformability compared to linear systems due to the breaking of superposition. It is well known that different types of nonlinearity yield vastly different effects on propagating signals, which raises the question of “what precise nonlinearity is the best for a given wave tailoring application?” Considering a one-dimensional spring-mass chain as a testbed, we couple the shape optimization of structures for tailored nonlinear constitutive responses with reduced-order nonlinear dynamical inverse design. Using minimization of peak kinetic energy transmission from impact as a case study, we identify ideal nonlinear constitutive responses and the geometries needed to achieve them. As part of this, we show the large sensitivity of this metric to small changes in nonlinearity, and thus the need for high precision, free-form nonlinearity tailoring. We validate our predictions using impact experiments in a chain of nonlinear springs and masses. This work sets the foundation for broader passive nonlinear mechanical wave tailoring material design, with applications to computing, signal processing, shock mitigation, and autonomous materials.
keywords: nonlinear dynamics, nonlinear mechanics, impact mitigation, nonlinear waves, inverse design
Introduction
@@
In this work, we introduce a method to create customizable wave tailoring materials via nonlinear bilevel inverse design. Namely, we optimize for the emergent dynamic response of a mesostructured material in the form of a one-dimensional (1D) spring-mass chain. To do this, we use a reduced order, discrete element model (DEM) simulation to identify an optimal nonlinear constitutive law for the given performance metric, and couple this to a unit-cell-scale, geometrically-nonlinear, free-form, shape optimization algorithm which designs a physical system that achieves the nonlinear constitutive property identified by the DEM (outlined in Fig. 1). We note that the 1D spring-mass chain is a prototypical nonlinear system that has classically formed the foundation for extensions into higher dimensions and continua [37, 34, 17], and, as part of our work herein, we illustrate extensions of our optimized unit-cells to two- and three-dimensional analogs (see Supplementary Information Note 1). Further, we demonstrate the clear need for high precision nonlinear response design, wherein the system dynamic performance is shown to be highly sensitive to the nonlinear coefficients of the constitutive law. Herein, we choose the maximum kinetic energy transmitted to the boundary of the system in response to an impact event, normalized by that of an otherwise-identical linear system, as our performance metric. The comparison between the linear and nonlinear response is particularly important, as it isolates the role of nonlinearity from other wave manipulating effects such as dispersion and dissipation. We demonstrate the full inverse design of this superior impact mitigation system, from identification of an ideal nonlinearity, to the design of the unit cell geometry, and experimental validation of the performance. An important feature to note, unlike some prior computational quasi-static nonlinear mechanical design strategies [22, 38, 24, 25], we do not use simplified or reduced order models for our underlying mesostructure design, which enables a broader design space and access to highly precise tailoring of nonlinear responses [23]. While we demonstrate our bilevel inverse design for the case of impact mitigation, we expect such an approach for the previously specified nonlinear wave transformation applications [1, 7, 8, 9, 10, 11, 12, 13] within the acoustic, phononic, and mechanical wave settings.
@@
Figure 1: Overview of the bilevel design flow. (A) Discrete element model (DEM) simulation of the entire system dynamics. (B) Identification of the optimal nonlinear constitutive law. (C) Shape optimization of a mesostructure to match the identified nonlinear constitutive law. (D) Mechanical and dynamic experimental system characterization.
Results
Reduced order discrete element model
@@ Results
The DEM (Fig. 1A) simulates the impact of a rigid “impactor” block onto a chain with $N$ unit cells and $a$ unit cell length. We represent the chain as lumped masses of mass $m$, interconnected by massless nonlinear springs and inter-site linear dampers in parallel, emulating the behavior of a viscoelastic material. The nonlinear spring consists of a linear spring behavior in tension and an up-to-third-order polynomial in compression, where the linear stiffness remains fixed at $c_{1}^{\ast}$ in both tension and compression (as shown in Fig. 1B). This results in non-dimensionalized equations of motion for our chain
@@
where dimensionless stiffnesses $c_{2}$ and $c_{3}$ are zero when that spring is in tension, $\zeta$ is the inter-site damping ratio, $x_{i}$ is the dimensionless displacement of the $i$th particle from its rest position, overdots represent the derivative with respect to nondimensional time, and all variables and parameters are normalized by combinations of $a$, $m$, and/or $c_{1}^{\ast}$, as is described in Supplementary Information Note 2. We note that while the masses are illustrated as $m$ in Fig. 1A, the nondimensional mass of each particle remains one as in Eq. (1). A variable mass and velocity rigid impactor is applied on the left, and a fixed boundary on the right. The model also incorporates a contact spring designed to facilitate the smooth contact and controlled release of the impactor during initial impact and rebound, respectively. The simulated dynamical response is acquired through the numerical integration of Eq. (1) via a Runge-Kutta algorithm. The non-dimensionalization of all variables and full details concerning the equations of motion are described in Supplementary Information Note 2.
Optimization of nonlinear constitutive law based on dynamical response
@@
Before searching for optimal nonlinear constitutive responses with our DEM, we set several bounds. First, we confine the unit cell strain to $1$ in compression, and set $c_{3}>0$ for simplicity. Second, we restrict our search range for nonlinear coefficients $c_{2}$ and $c_{3}$ to ensure positive strain energy throughout the entire compression range. We note that keeping the linear stiffness constant, we exclude essential nonlinearities [14]. By examining the polynomial’s properties within this range, we classify the quasi-static response into three distinct zones, “bistability”, “monotonic increase”, and “local maximum”, as shown in Fig. 2A. Bistability (magenta area) denotes the existence of both a local maximum and minimum other than the boundaries (the local minimum does not need to fall below zero). Monotonic increase (blue area) denotes the absence of extrema. Local maximum (green area) signifies the presence of a local maximum (no local minimum existed) within the range of the length of one unit cell. More details concerning these zones is given in Supplementary Information Note 3.
@@
Figure 2: Identification of optimal nonlinear constitutive response via DEM simulation for a single impact condition. (A) Feasible solutions of nonlinear spring coefficients $c_{2}$ and $c_{3}$. The black line represents $c_{2}=-\sqrt{3c_{3}}$, the red line indicates $c_{2}=-(1+3c_{3})/2$, the blue line is $c_{2}=-3c_{3}$, and the green line is the zero strain energy throughout the whole range, $c_{2}=-3/2-3c_{3}/4$. (B) Non-dimensional force-extension relationship of the best performing nonlinear spring. (C) Ratio of maximum kinetic energy at the last particle in the nonlinear chain to the initial impact energy as a function of nonlinear spring coefficients for the impact condition of $M/M_{0}=0.05$ and $V/V_{0}=1$, with $\zeta=0.01$. The lines from (A) are overlaid, and the star marker denotes the point of best performance. Normalized kinetic energy of the (D) best performing nonlinear and (E) linear material.
@@
Within these bounds, we first optimize for a singular impact condition ($M/M_{0}=0.05$ and $V/V_{0}=1$), where $M_{0}$ is half the mass of the chain, $V_{0}$ is the linear sound speed, and $M$ and $V$ are the dimensional impactor mass and velocity, respectively. The material is composed of $20$ particles and $\zeta=0.01$. We vary the nonlinear coefficients of the springs $c_{2}$ and $c_{3}$, aiming to minimize the maximum kinetic energy experienced at the end of the material ($KE_{non}$) normalized by that of a linear system ($KE_{lin}$) which has all of the same properties except $c_{2}=c_{3}=0$. Given dimensional particle displacement from its rest position $x_{i}^{\ast}$, kinetic energy of the $i$th particle is defined as $m(dx_{i}^{\ast}/dt^{\ast})^{2}/2$, where $t^{\ast}$ is dimensional time, such that the analogous dimensionless kinetic energy is $\dot{x_{i}}^{2}/2$. Figure 2C quantifies the ratio of maximum kinetic energy at the end of the material to the initial impact energy ($TE$), as a function of the coefficients of the nonlinear spring. The best performance in Fig. 2C corresponds to nondimensional spring force $f(\Delta x)=\Delta x+5.81\Delta x^{2}+7.08\Delta x^{3}$, where $\Delta x$ is the spring extension and positive $f$ denotes tension. This optimal spring is plotted in Fig. 2B, and is denoted by the star marker in the inset of Fig. 2C. A point of particular note about this identified optimum, is that it is not of the form one would expect based on the conventional design approach for bistable energy absorption at lower rates [39], where net positive energy is locked into strain energy when snapping from the undeformed state to its second stable equilibrium (i.e., the area under the curve from $\Delta x=0$ to the unstable equilibrium point (at $\Delta x\approx-0.2$), is less than the area under the curve from the unstable equilibrium to the second stable equilibrium point (at $\Delta x\approx-0.6$). Rather in the case of our optimum, the first area is smaller than the second. Around this point of best performance, significant performance sensitivity to the constituent nonlinear parameters can be seen (see the zoomed-in view in Fig. 2C). To describe this sensitivity more quantitatively, in Supplementary Information Note 4 we calculate the gradient of the inset in Fig. 2C, which shows $\partial(\log_{10}(KE_{non}/TE))/\partial c_{2}$ and $\partial(\log_{10}(KE_{non}/TE))/\partial c_{3}$ approaching ±$\pm 10$. This metric thus suggests that, e.g., a $0.1$ difference in $c_{2}$ or $c_{3}$ can result in up to a 10×10$\times$10 × change in peak $KE$. However, we emphasize that near our chosen optimum, the sensitivity is significantly lower, with $\partial(\log_{10}(KE_{non}/TE))/\partial c_{2}=0.32$ and $\partial(\log_{10}(KE_{non}/TE))/\partial c_{3}=0.07$. Spatiotemporal responses of kinetic energy of the optimal nonlinear and linear chain are shown in Fig.  2(D) and (E), respectively. In Fig. 2(D) and (E), while both materials exhibit an initially sharp pulse, energy from the optimal nonlinear material is trapped around particle $15$, preventing further transmission.
@@
In the subsequent analysis, we conduct an optimization wherein we look for optimal nonlinear coefficients for varied impactor mass and velocities. We use the same ratio of the maximum $KE$ of the last particle of the nonlinear material normalized by that of the linear material as a performance metric (denoted henceforth as “$KE$ ratio”), wherein a lower ratio indicates superior performance of the nonlinear material. The optimal (minimum) $KE$ ratios with respect to $M/M_{0}$ and $V/V_{0}$ and corresponding nonlinear spring parameters $c_{2}$ and $c_{3}$ can be seen in Fig. 3. In contrast to Fig. 2, we use lower damping ($\zeta=0.005$), chosen to emulate that of the polycarbonate springs used in our experimental realization. Additional simulation results of $KE$ ratios for increased damping and greater discreteness (more unit cells) are available in Supplementary Information Note 5 and indicate the potential for $KE\leavevmode\nobreak\ \text{ratio}<10^{-2}$ in the latter case. The damping value used in Fig. 3 was chosen by measuring the low-amplitude resonance of a single connector and nonlinear spring unit (see Supplementary Information Note 6). This is an important point to highlight, in that the characterizing the damping at low amplitudes allows us to temporarily discard its interplay with the nonlinear spring. This damping can be thought to stem from the intrinsic damping of the polycarbonate. Using the model of the nonlinear spring with the linear damper in parallel, thus allows us to simulate greater effective damping that may come from their dynamical interplay at higher amplitudes. Simulation cases of the nonlinear material where self-contact occurs are discarded from consideration of the optimal performance.
@@
Figure 3: Identification of optimal nonlinear coefficients via DEM simulation for varied impactor conditions. (A) Optimal kinetic energy ratio ($KE$ ratio) as a function of impact conditions and corresponding nonlinear spring coefficients (B) $c_{2}$ and (C) $c_{3}$ for a material of $N=20$ and $\zeta=0.005$. The color bar in (A) is saturated at a $KE$ ratio of unity. The dashed black line denotes the onset of self contact ($\Delta x<-1$) in the linear system. The dot in (A) denotes the high performing case used for mesostructure design.
@@
There is a clear boundary where the nonlinear chain does not outperform the linear, which is correlated with the occurrence of self-contact within the linear chain (dashed black line in Fig. 3). At impactor velocities and masses below this threshold, the nonlinear materials exhibit significantly enhanced mitigation effectiveness. As can be seen in Fig. 3A, at a relatively low to moderate level of impactor mass ($\sim 0.01$ to $0.2M_{0}$), optimal nonlinear materials exhibit a mitigation capability ($KE$ ratio) that is over ten times better than the linear material. As the impactor mass and velocity increase and cross over the dashed black line (orange to red area near the dashed line in Fig. 3A), the priority shifts to preventing contact between unit cells, leading to a comparatively impaired energy-absorbing performance. In future studies, such self contact could be explored as a form of nonlinearity, and a design feature instead of a constraint.
@@
A specific optimal solution (pair of coefficient ratios) was chosen from a set of multiple solutions showing high performance $KE$ ratio. We further downselected, looking for a solution also with relatively low strain, making it more amenable to experimental implementation. The chosen solution is a nondimensional nonlinear mechanical response of the form $f(\Delta x)=\Delta x+87\Delta x^{2}+1778\Delta x^{3}$ in compression, with an impact condition of $M/M_{0}=0.01$ and $V/V_{0}=0.22$ (marked in Fig. 3A). This nonlinear constitutive law is used as the target for the mesostructure design, with the goal of finding a geometry that yields that nonlinear mechanical response.
Shape optimization for desired nonlinear constitutive law
@@
In order to find a spring geometry that gives the desired nonlinear response, a two-dimensional shape optimization approach is taken, using a level-set optimization method (as detailed further in the Methods and Ref. [23]). In summary, a third order polynomial is fit to the calculated force-displacement behavior of the structure, and the objective of the optimization problem is taken as the ratio of nonlinear to linear terms (e.g., as shown in Fig. 1C and Fig. 4A). The exact form of the objective function is
@@
where $c_{i}$ and $c_{j}$ represent the current polynomial coefficients, and $T_{i}$ and $T_{j}$ represent the target coefficients. By taking the ratio of the polynomial terms, the nonlinearity of the structure is decoupled from the linear stiffness, allowing the optimizer more design freedom, leading to more robust convergence. To further aid in the navigation of the design space, we select an initial condition for our optimization process, which displays qualitatively similar behavior to that desired (see Supplementary Information note 6).
@@
The boundary conditions applied are depicted in Fig. 4B (fixed on top, roller on bottom, applied displacement on bottom 10% of the right boundary). When realizing our nonlinear chain, we employ comparatively rigid frames around the designed spring to mimic fixed boundary conditions (as can be seen in Fig. 1D and Fig. 4C) and rigid connectors between the springs to allow relative movement of the masses. With the addition of these components we have a unit cell length, $a$, and a nonlinear spring design domain length that encompasses only a portion of this larger unit cell. We call this portion of $a$ the spring length, and denote it by $a_{s}$ (and, hereafter, the subscript $s$ is used to refer to parameters defined on the scale of the spring). Figure 4E highlights the difference between these two length scales. Because the DEM-identified polynomial constitutive law is expressed as a function of strain, and the strain experienced by the spring across $a_{s}$ is different than that experienced by the entire unit cell across $a$ for a fixed applied displacement, the targeted polynomial is therefore scaled accordingly. We take the ratio of nonlinear terms as $R=(c_{i}\epsilon^{i})/(c_{1}\epsilon)$ and $R_{s}=(c_{s,i}\epsilon_{s}^{i})/(c_{s,1}\epsilon_{s})$ on the scales of $a$ and $a_{s}$, respectively, with $i$ representing the order (or power) of the term and $\epsilon$ representing the maximum strain experienced on the corresponding length scale. In order for equivalent degrees of nonlinearity to be displayed at different scales, we take $R=R_{s}$ at the maximum strain considered in each scale, and solve for updated $c_{i}$ or $c_{s,i}$ terms.
@@
The optimization target (recall, on the unit cell scale, identified above as $f(\Delta x)=\Delta x-87\Delta x^{2}+1778\Delta x^{3}$) can therefore be expressed on the spring length scale as $f(\Delta x_{s})=\Delta x_{s}-41.064\Delta x_{s}^{2}+396.11\Delta x_{s}^{3}$, by setting $R=R_{s}$ and solving for $c_{s,i}$. The final optimized structure, shown in Fig. 4A-B, achieved a simulation force-displacement law on the spring scale of $f(\Delta x)=\Delta x_{s}-40.712\Delta x_{s}^{2}+397.535\Delta x_{s}^{3}$, and an experimental one of $f(\Delta x)=\Delta x_{s}-40.654\Delta x_{s}^{2}+397.274\Delta x_{s}^{3}$ (shown in Fig. 4D), resulting in an experimental percent difference between targeted and obtained polynomial ratios of $0.294\%$ for the third order ratio and $0.998\%$ for the second order ratio.
@@
Figure 4: Shape optimization to identify geometry giving DEM identified desired constitutive law, and experimental realization of the spring and chain. A-B) Results of the shape optimization, with the target and final force-displacement curves shown in (A), and the optimized design shown in (B). C) The fabricated polycarbonate unit cell, consisting of four optimized springs (as seen in (B)) and a rigid frame to impose boundary conditions. D) The quasi-static test of the unit cell shown in (C), compared against the behavior of a single spring with perfectly imposed boundary conditions as simulated via COMSOL FEM simulation. E) The full chain of 20 unit cells, hung from a frame. The chain is clamped to the left of the leftmost unit cell, imposing a zero displacement boundary condition. The impact occurs at the right end of the chain. The unit cell and spring length scales are labeled.
@@
As the multilevel optimization methodology is a loop, in which inputs from the DEM and shape optimization feed back into one another, the design process can be followed in either direction, which can be exploited for exploration of the space. For example, while one may begin with a desired impact condition, which is then fed into the DEM simulation followed by the mesostructure optimization, one might just as easily begin with some nonlinearity and feed this into the DEM simulation to explore what impact conditions it might perform well (or poorly) for. Similarly, one might find that the physical design domain requires constraints, or that optimized structures exhibit structural damping, which diverges significantly from predictions, and these updated considerations can be flowed back up to the DEM simulation to search for the updated optimum conditions.